# LoRA-Mixer: Coordinate Modular LoRA Experts Through Serial Attention Routing

### v1

@@ Abstract
Recent efforts to combine low-rank adaptation (LoRA) with mixture-of-experts (MoE) for adapting large language models (LLMs) to multiple tasks, yet exhibit prevailing limitations: they either swap entire attention/feed-forward layers for switch experts or bolt on parallel expert branches, diluting parameter efficiency and task fidelity. We propose the LoRA-Mixer, a modular and lightweight MoE framework that integrates LoRA experts. Our core innovation lies in replacing the projection matrices of the attention module’s input/output linear layers with dynamically routed, task-specific LoRA experts. This design ensures seamless compatibility with diverse foundation models—including transformers and state space models (SSMs)—by leveraging their inherent linear projection structures. The framework supports two operational paradigms: (1) joint optimization of LoRA experts and routing mechanisms via a novel hard-soft routing strategy, or (2) direct deployment of pre-trained, frozen LoRA modules sourced from external repositories. To enable robust router training with limited data while ensuring stable routing decisions and maximizing expert reuse, we introduce an adaptive Specialization Balance Loss (SBL) that jointly optimizes expert balance and task-specific alignment. Extensive experiments on seven benchmark datasets, including MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval, demonstrate the effectiveness of LoRA-Mixer. On datasets such as GSM8K, HumanEval, and MedQA, LoRA-Mixer achieves significant improvements of 7.61%, 4.88%, and 3.08% over the base models, respectively. Compared with the state-of-the-art methods, LoRA-Mixer achieves additional improvements of 1.09%, 1.45%, and 1.68%, respectively, using only 48% of the parameters, demonstrating its efficiency and strong performance.
1 Introduction
@@
Large Language Models (LLMs) have achieved unprecedented proficiency in general-purpose reasoning and generation, yet their adaptation to specialized downstream domains remains computationally prohibitive, requiring significant resources for full-scale fine-tuning [1, 2]. To mitigate these resource demands, parameter-efficient fine-tuning (PEFT) methods [37, 38, 39, 40, 41, 42] have emerged as a scalable paradigm for task-specific adaptation. Among these, Low-Rank Adaptation (LoRA) [4, 5] has demonstrated particular efficacy, operating through low-rank decomposition of updates to the pretrained weights—enabling efficient tuning with minimal parameter overhead. Recent work has explored modularly composing independently trained LoRA modules as a promising strategy for multitask adaptation; however, naive composition can result in interference between task-specific subspaces, limiting their synergistic potential[15, 14]. This limitation has motivated exploration of mixture-of-experts (MoE) architecture [3, 6], which treat each task-specific LoRA as an expert and sparsely activate and fuse the experts. Recent studies demonstrate promising directions in hybrid Lo-
@@
Figure 1: MoE and LoRA-MoE integration methods. (1) and (2): replace the attention or feedforward layers with switch experts; (3) and (4): introduce LoRA experts branches in parallel with the attention or feedforward layers, and fuse the output into the main branch. Our LoRA-Mixer (right) applies mixture of LoRA experts to the projection layers, which can effectively leverages the attention mechanism.
RA-MoE frameworks [14, 13, 15, 16, 35, 34, 23, 24], aiming to enhance model performance on complex tasks across multi-domain datasets while preserving the parameter efficiency of fine-tuning.
The central challenge in composing multiple pre-trained LoRAs lies in effectively synergizing them—achieving enhanced performance across constituent tasks while minimizing training overhead and preserving their distinct, task-specific characteristics. Conventional methods integrating LoRA into mixture-of-experts (MoE) architectures typically follow one of two paradigms: (1) directly substituting attention or feedforward layers with LoRA-based switchable experts [13, 23], emulating the classical MoE structure [6, 3]; or (2) introducing parallel branches of LoRA experts whose outputs are subsequently fused into the primary model pipeline [14, 15], as illustrated in Fig.1. While these strategies have demonstrated promising empirical results [14, 35, 52, 53, 54, 55], they still encounter fundamental limitations. Specifically, the vanilla MoE-inspired paradigm necessitates joint training across all expert modules, significantly increasing training data demands and restricting the modular reuse and transferability of pre-trained LoRAs. Conversely, the parallel LoRA-expert approach circumvents the inherent attention or state transition mechanisms, resulting in simplistic output fusion and suboptimal overall integration. Additionally, auxiliary losses commonly employed for routing optimization inherently promote uniform load balancing among experts, diminishing the capacity for nuanced task-awareness [13]. These limitations motivate the development of a more flexible, plug-and-play framework that is model-architecture agnostic, compatible with both Transformer and state-space model (SSM) architectures [7], capable of training an efficient and discriminative routing mechanism with minimal computational and data demands, and maximizing the reuse and transferability of independently pre-trained LoRA modules.
@@
In this paper, we introduce LoRA-Mixer, a novel framework designed to efficiently synergize multiple pre-trained LoRA modules by treating them as dynamic, pluggable memory cells. LoRA-Mixer equips the linear projection layers of the original model with mixed LoRA experts, enabling these experts to directly leverage the effectiveness of the core attention or state-transition mechanisms. LoRA-Mixer supports LoRA modules sourced from external repositories, independently trained, or jointly trained through hard routing strategies, allowing seamless plug-and-play usage across various tasks and domains. Importantly, our method significantly reduces the necessity for training data or extensive re-adaptation, requiring only minimal additional data to effectively train the routing mechanism. Consequently, LoRA-Mixer is particularly suitable for constructing large-scale, modular language models characterized by task-specific memory, computational efficiency, and strong transferability. To further enhance efficiency and maintain routing effectiveness, we propose a novel Router Specialization Balancing Loss (RSBL). RSL aligns routing decisions with token-level expert usage, maintaining moderate entropy to encourage exploratory behavior. During inference, we employ sparse top-$K$ fusion, effectively balancing computational cost and scalability without compromising expert selectivity. Extensive evaluations conducted on seven benchmark datasets—MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval—demonstrate that integrating LoRA-Mixer substantially improves model performance across all evaluated tasks. Additionally, cross-domain experiments confirm the versatility and adaptability of our proposed framework. Compared to state-
@@
of-the-art approaches, LoRA-Mixer achieves significant efficiency, using only 48% of the training parameters while improving performance by 1.09%, 1.45%, 1.19%, and 0.85% on GSM8K, HumanEval, SST-2, and CoLA benchmarks, respectively.
2 Related Work
@@
PEFT For LLMs Low-rank adaptation (LoRA) [12, 11, 4, 10, 47, 49, 50, 51] effectively fine-tunes large models by learning a low-rank matrix and freezing the original weights. While effective for a single task, its task-specific nature limits generalization. Recent work combines LoRA with mixture of experts (MoE) [13, 14, 25, 15, 16] to achieve dynamic adaptation. For example, MixLoRA [13] uses LoRA experts for top-k routing in FFN, improving multi-task performance but suffering from gradient entanglement issues. MoLE [14] combines LoRA layers via gating but lacks sparse routing. LoraHub [15] performs gradient-free few-shot combination of LoRA modules for unseen tasks but struggles with complex semantics due to lack of gradient optimization and dynamic routing. Other methods [18, 19, 45, 46, 48, 60] explore flexible routing mechanisms to improve the model’s adaptability. However, these methods usually introduce additional routing networks or optimization targets, resulting in instability during training, limiting their application in actual multi-task or low-resource scenarios.
@@
Mixture of Experts In recent years, the mixture of experts (MoE) architecture has attracted much attention as a promising LLM expansion paradigm. By selectively activating a subset of expert modules for each input, MoE allows the model to scale capacity without linearly increasing the amount of computation. As a result, more and more large models have adopted MoE, including GLaM[20], Switch Transformers[6], and the recent DeepSeek series[21, 22]. These advances indicate that MoE is becoming a mainstream architectural trend in the development of next-generation base models. Among them, GLaM[20] and Switch Transformer[6] build a mixture of experts (MoE) model in the FFN module, and use a sparse activation mixed expert architecture to expand the model capacity and achieve better performance. LLaVA-MoLE[23] uses a top-1 strategy to route tokens to domain-specific expert models, thereby alleviating data conflicts and achieving continuous performance improvements over the ordinary LoRA baseline. LoRAMoE[24] uses routers to integrate LoRA experts while retaining general knowledge. HMoRA[35] combines the layered fine-tuning methods of MoE and LoRA, and gradually switches the routing strategy as the number of layers increases. MoLA[34] assigns different numbers of experts at different levels, proving that deeper layers often require more experts. In our method, we use mixed LoRA experts in the projection layer and optimize the load loss. LoRA-Mixer is a more fine-grained MoE construction method that is independent of the architecture and can simultaneously maintain expert expertise, computational sparsity, and routing adaptability. In addition, LoRA-Mixer can reuse existing LoRA modules and only requires very little data to train routing. While saving computing resources, it achieves the expansion of model capacity and the improvement of generalization ability.
3 Method
@@
In this section, we introduce LoRA-Mixer, a flexible and pluggable mixture of experts (MoE) framework for combing multiple LoRA experts of LLMs.
@@ 3.1 Preliminaries
LoRA [4] is a parameter-efficient fine-tuning method that adapts large pre-trained models by introducing low-rank updates to the original weight matrices, instead of updating them directly. Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, LoRA freezes $W_{0}$ and introduces a trainable low-rank decomposition $\Delta W=BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}\,,\ B\in\mathbb{R}^{d_{\text{out}}\times r}\,,\ r\ll\min(d_{\text{in}},d_{\text{out}})$. The adapted transformation becomes:
$y=(W_{0}+BA)\mathbf{x}=W_{0}x+B(A\mathbf{x}),$ (1)
where $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ is the input token representation. This technique significantly reduces trainable parameters to $r(d_{\text{in}}+d_{\text{out}})$, enabling scalable fine-tuning with limited resources.
Mixture-of-Experts [3] is a sparse neural architecture where each input token is processed by a small subset of expert networks. Given $K$ experts and a router that produces a score vector $G(\mathbf{x})\in\mathbb{R}^{K}$, a softmax is applied to obtain the routing distribution:
@@ 3.2 LoRA-Mixer for Compositing LoRAs
LoRA Experts Acquirement. Our proposed LoRA-Mixer framework is highly flexible and supports the integration of LoRA modules from diverse sources. In one common scenario, users may download pre-trained LoRA adapters from public repositories such as LoRAHub [15], which currently hosts 196 high-quality LoRA modules across a wide range of domains. These can be directly composed using LoRA-Mixer with minimal additional data. Alternatively, users may independently train domain-specific LoRA modules tailored to their own datasets. For scenarios requiring joint training of multiple LoRA modules on a heterogeneous, labeled dataset, LoRA-Mixer further supports a hard-routing strategy. Specifically, we fix the routing module and apply a deterministic routing scheme based on known domain labels. Given a domain ID $d\in\{1,\dots,K\}$ associated with each training instance, all tokens within the sample are routed exclusively to expert $d$. This design enables efficient joint optimization while maintaining expert modularity. Collectively, these capabilities make LoRA-Mixer a versatile and scalable framework for composing heterogeneous LoRA modules. The overall architecture is illustrated in Figure 2.
@@ 3.3 Specialization Balance Loss for Routing Optimization
The next step is to optimize the expert router. Although previous studies  [22, 21, 13] introduced auxiliary losses to align the average gating score with the expert utilization to promote load balancing, we observed that this approach overemphasized consistency, resulting in an overly balanced distribution of experts. In this case, all experts are forced to be used equally regardless of the input semantics. This hinders effective routing and often requires more training data. To ensure that the expert load is balanced and the routing is input-aware, we propose an improved optimization objective called Route-Specialization Balance Loss (RSL).
We introduce a selectivity-aware regularization term that regulates the entropy of routing distribution to enhance the auxiliary loss, guiding routers to make more discriminative expert choices instead of blindly averaging activations. Formally, let $\bar{p}_{i}$ denote the average soft route score (across tokens) and $\bar{f}_{i}$ denote the normalized score of the token assigned to expert $i$ in the first $k$ routes. The RSL loss function is defined as:
@@
Where $\alpha$ controls the strength of the balanced consistency term, $\lambda$ is a small positive coefficient of the entropy regularizer:
@@
which is calculated based on the routing distribution $p(\mathbf{x})$ of each token. The first term enforces that routing intentions are consistent with actual usage, while the second term penalizes excessively high entropy in expert selection, thereby promoting more specialized expert assignments.
This combination ensures that the router not only maintains overall balance, but also forms a preference structure that avoids the problem that every expert is used equally regardless of the input semantics. We have analyzed RSL and traditional auxiliary loss from a theoretical perspective. Please refer to Appendix C for detailed derivation. For the balance loss in training, please refer to the Appendix  D.
Routing Optimization. After we prepare all LoRAs, we apply a soft training process on router. To prevent the expert knowledge in the first phase from being contaminated, we introduce a regularization term to penalize deviations from the previously learned expert parameters. Let the first phase parameters of expert $i$ be $\theta_{i}^{(0)}$ and the current parameters be $\theta_{i}$. We define the regularization term as:
@@
where $\mathcal{C}$ represents the set of constrained experts and $\beta$ controls the regularization strength. This regularization term constrains the sensitive experts to stay close to their original knowledge while still allowing other experts to adjust. For complex tasks, we support expert-level control, thus enabling flexible multi-expert learning.
To ensure that all experts can obtain meaningful gradients during the joint training process and promote stable optimization of the routing balance loss, we adopt soft expert fusion in training. Specifically, the router outputs the softmax routing scores $\mathbf{p}_{b,t}\in\mathbb{R}^{K}$ of all experts and fuses them, thereby achieving a fully differentiable mixture of LoRA experts. Although soft routing provides stable optimization and gradient propagation for all experts, its combination with the auxiliary loss leads to the problem of extremely balanced expert usage, that is, all experts are activated to the same extent regardless of the input semantics. To address this limitation, we introduce the Route-Specialization Balancing Loss (RSL) 3.3, which promotes the specialization of experts without sacrificing load balancing. The total loss during joint training becomes:
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{task}}}_{\text{Supervised Objective}}+\underbrace{\alpha\cdot\mathcal{L}_{\text{RSL}}}_{\text{Routing Specialization}}+\underbrace{\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2}}_{\text{Expert Preservation}},$ (8)
where $\mathcal{L}_{\text{task}}$ denotes the standard task loss (e.g., Cross-Entropy Loss). $\alpha$ is weighting factor for RSL. $\mathcal{L}_{\text{preserve}}$ is the expert regularization loss described above, scaled by $\beta$.
At inference time, we adopt a Top-3 routing strategy. By separating soft fusion at training time from sparse activation at inference time, we can ensure that the model obtains robust gradient propagation during routing optimization while performing efficient inference.
4 Experiment
To evaluate the effectiveness of our proposed LoRA-Mixer framework, we conduct experiments with LoRAs finetuned in five domains: Medical QA, Commonsense Reasoning, Natural Language Understanding, Mathematical Reasoning, and Coding Ability.

### v2

@@ Abstract
Recent attempts to combine low-rank adaptation (LoRA) with mixture-of-experts (MoE) for multi-task adaptation of Large Language Models (LLMs) often replace whole attention/FFN layers with switch experts or append parallel expert branches, undermining parameter efficiency and limiting task specialization. We introduce LoRA-Mixer, a modular MoE framework that routes task-specific LoRA experts into the core projection matrices of the attention module (input/output linear layers), rather than primarily targeting FFN blocks. The design delivers fine-grained token-level specialization by fully exploiting the attention mechanism, while remaining drop-in compatible with Transformers and state-space models (SSMs) as the linear projection layers are ubiquitous. To train robust routers from limited data while promoting stable, selective decisions and high expert reuse, LoRA-Mixer employs an adaptive Routing Specialization Loss (RSL) that jointly enforces global load balance and input-aware specialization via an entropy-shaping objective. The framework supports two regimes: (i) joint optimization of adapters and router with a differentiable hard–soft top-$k$ routing scheme, and (ii) plug-and-play routing over frozen, pre-trained LoRA modules sourced from public repositories. Across 15 benchmarks—including MedQA, GSM8K, HumanEval, and GLUE—RSL-optimized LoRA-Mixer outperforms state-of-the-art routing and LoRA-MoE baselines while using 48% of their trainable parameters, with gains of +3.79%, +2.90%, and +3.95% on GSM8K, CoLA, and ARC-C, respectively. Cross-model transfer and adapter reuse experiments further demonstrate the approach’s versatility and data efficiency. Our code can be obtained at https://github.com/hustcselwb/LoRA-Mixer.
1 Introduction.
@@
Large Language Models (LLMs) have achieved unprecedented proficiency in general-purpose reasoning and generation, yet their adaptation to specialized downstream domains remains computationally prohibitive for full-scale fine-tuning Brown et al. (2020); Touvron et al. (2023). To mitigate these resource demands, parameter-efficient fine-tuning (PEFT) methods Li & Liang (2021); Zaken et al. (2021); Liu et al. (2022); Houlsby et al. (2019); Liu et al. (2021) have emerged as a scalable paradigm. Among them, Low-Rank Adaptation (LoRA) Hu et al. (2022); Tian et al. (2024) has demonstrated particular efficacy, operating through low-rank decomposition of updates to the pretrained weights—enabling efficient tuning with minimal parameter overhead. Recent work has explored modularly composing independently trained LoRA modules as a promising strategy for multitask adaptation; however, naive composition can result in interference between task-specific subspaces, limiting their synergistic potential Huang et al. (2023); Wu et al. (2024). This limitation has motivated exploration of mixture-of-experts (MoE) architecture Shazeer et al. (2017); Fedus et al. (2022), which treat each task-specific LoRA as an expert and sparsely activate and fuse the experts. Recent studies demonstrate promising directions in hybrid LoRA-MoE frameworks Wu et al. (2024); Li et al. (2024a); Huang et al. (2023); Zhao et al. (2024); Liao et al. (2025); Gao et al. (2024); Chen et al. (2024); Dou et al. (2023), aiming to enhance model performance on complex tasks across multi-domain datasets while preserving the parameter efficiency of fine-tuning.
@@
Figure 1: LoRA and MoE integration methods. (1) and (2): replace the attention or FNNs with switch experts; (3) and (4): introduce LoRA experts branches in parallel with the attention or feedforward layers, and fuse the output into the main branch. Our LoRA-Mixer (right) applies mixture of LoRA experts to the projection layers, which can effectively leverages the attention mechanism.
The central challenge in composing pre-trained LoRAs is to realize synergy—improving performance across constituent tasks—without inflating training cost or erasing task-specific inductive biases. Existing LoRA–MoE integrations largely follow two patterns: (i) replacing attention or FFN blocks with LoRA-based switch experts in a standard MoE setup Li et al. (2024a); Chen et al. (2024); Fedus et al. (2022); Shazeer et al. (2017); or (ii) attaching parallel LoRA branches whose outputs are fused back into the main path Wu et al. (2024); Huang et al. (2023) (Fig. 1). Despite encouraging results Wu et al. (2024); Liao et al. (2025); Krajewski et al. (2024); Zhang et al. (2019); Li et al. (2022); Gross et al. (2017), switch-style designs typically require joint training of all experts, increasing data demands and hindering modular reuse and transfer of off-the-shelf LoRAs. Parallel-branch schemes sidestep native attention or state-transition pathways, yielding shallow output fusion and weak integration. Moreover, common auxiliary routing losses emphasize uniform load balancing, suppressing input- and task-aware specialization Li et al. (2024a; 2026); Zhang et al. (2026c). These issues motivate a plug-and-play, architecture-agnostic alternative—compatible with both Transformers and state-space models (SSMs) Gu & Dao (2023); Zhang et al. (2026a; b)—that learns discriminative routing with minimal compute and data while maximizing reuse of independently trained LoRA modules.
@@
In response, we introduce LoRA-Mixer, a novel framework designed to efficiently synergize multiple pre-trained LoRA modules by treating them as dynamic, pluggable memory cells. LoRA-Mixer equips the linear projection layers of the original model with mixed LoRA experts, enabling these experts to directly leverage the core attention or state-transition mechanisms. To further enhance efficiency and maintain routing effectiveness, LoRA-Mixer adopts a novel Router Specialization Balancing Loss (RSL) to align routing decisions with token-level expert usage, maintaining moderate entropy to encourage exploratory behavior. During inference, we employ sparse top-$K$ fusion, effectively balancing computational cost and scalability without compromising expert selectivity. LoRA-Mixer supports LoRA modules sourced from external repositories, independently trained, or jointly trained through hard routing strategies, allowing seamless plug-and-play usage across various tasks and domains. Importantly, our method significantly reduces the necessity for training data or extensive re-adaptation, requiring only minimal additional data to effectively train the routing mechanism. Consequently, LoRA-Mixer is particularly suitable for constructing large-scale, modular language models characterized by task-specific memory, computational efficiency, and strong transferability. Extensive evaluation on 15 benchmark datasets (MedQA, GLUE, GSM8K, ARC-E, ARC-C, HumanEval, PIQA, HellaSwag, and BoolQ) demonstrates that integrating LoRA-Mixer significantly improves model performance across all evaluated tasks. With only 48% of the parameters of existing methods, the RSL-optimized LoRA-Mixer outperforms the state-of-the-art routing and LoRA-MoE baseline models, achieving improvements of +3.79%, +2.90%, and +3.95% on GSM8K, CoLA, and ARC-C, respectively. Cross-model transfer and adaptive module reuse experiments further demonstrate the method’s versatility and data efficiency.
2 Related Work.
PEFT For LLMs Low-rank adaptation (LoRA) Hayou et al. (2024); Zhou et al. (2024); Hu et al. (2022); Xu et al. (2024); Sheng et al. (2023); Biderman et al. (2024); Li et al. (2023); Zeng et al. (2025); Lin et al. (2025)
@@
effectively fine-tunes large models by learning a low-rank matrix and freezing the original weights. While effective for a single task, its task-specific nature limits generalization. Recent work combines LoRA with mixture of experts (MoE) Li et al. (2024a); Wu et al. (2024); Ouyang et al. (2025); Huang et al. (2023); Zhao et al. (2024); Song et al. (2022; 2023); Hu et al. (2025) to achieve dynamic adaptation. For example, MixLoRA Li et al. (2024a) uses LoRA experts for top-k routing in FFN, improving multi-task performance but suffering from gradient entanglement issues. MoLE Wu et al. (2024) combines LoRA layers via gating but lacks sparse routing. LoraHub Huang et al. (2023) performs gradient-free few-shot combination of LoRA modules for unseen tasks but struggles with complex semantics due to lack of gradient optimization and dynamic routing. Other methods Li et al. (2025); Huang et al. (2025); Yang et al. (2024); Chen et al. (2023); Luo et al. (2024); Li et al. (2024b); cite11 explore flexible routing mechanisms to improve the model’s adaptability. However, these methods usually introduce additional routing networks or optimization targets, resulting in instability during training, limiting their application in actual multi-task or low-resource scenarios.
@@
Mixture of Experts In recent years, the mixture of experts (MoE) architecture has attracted much attention as a promising LLM expansion paradigm. By selectively activating a subset of expert modules for each input, MoE allows the model to scale capacity without linearly increasing the amount of computation. As a result, more and more large models have adopted MoE, including GLaMDu et al. (2022), Switch TransformersFedus et al. (2022); Song et al. (2025); Zhang et al. (2025a; b); Ye et al. (2025), and the recent DeepSeek seriesLiu et al. (2024); Guo et al. (2025a). These advances indicate that MoE is becoming a mainstream architectural trend in the development of next-generation base models. Among them, GLaMDu et al. (2022) and Switch TransformerFedus et al. (2022) build a mixture of experts (MoE) model in the FFN module, and use a sparse activation mixed expert architecture to expand the model capacity and achieve better performance. LLaVA-MoLEChen et al. (2024) uses a top-1 strategy to route tokens to domain-specific expert models, thereby alleviating data conflicts and achieving continuous performance improvements over the ordinary LoRA baseline. LoRAMoEDou et al. (2023) uses routers to integrate LoRA experts while retaining general knowledge. HMoRALiao et al. (2025) combines the layered fine-tuning methods of MoE and LoRA, and gradually switches the routing strategy as the number of layers increases. MoLAGao et al. (2024) assigns different numbers of experts at different levels, proving that deeper layers often require more experts. Notably, we only utilize hybrid LoRA experts in the core projection layer, resulting in extremely high parameter efficiency. By applying RSL to LoRA-Mixer, we can efficiently reuse LoRA modules and require minimal training data, saving computational resources while expanding model capacity and improving generalization.
3 Method.
@@
In this section, we introduce LoRA-Mixer, a flexible and pluggable Mixture of Experts (MoE) framework for combining LoRA experts from multiple LLMs.
@@ 3.1 Preliminaries
Mixture-of-Experts is a sparse neural architecture where each input token is processed by a small subset of expert networks. Given $K$ experts and a router that produces a score vector $G(\mathbf{x})\in\mathbb{R}^{K}$, a softmax is applied to obtain the routing distribution:
@@
Auxiliary loss. To encourage balanced expert utilization, MoE training introduces an auxiliary loss. We define the expected routing probability and top-1 usage as
$\bar{p}_{i}=\mathbb{E}_{x\sim\mathcal{D}}[p_{i}(x)],\qquad\bar{f}_{i}=\mathbb{E}_{x\sim\mathcal{D}}\big[\mathbb{I}(i=\arg\max_{j}p_{j}(x))\big],\qquad L_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\,\bar{f}_{i}.$ (3)
where $\bar{p}_{i}$ denotes the average routing intention and $\bar{f}_{i}$ the empirical usage. We demonstrate in Appendix A.17 why the auxiliary loss leads to over-averaging.
@@ 3.2 LoRA-Mixer for Compositing LoRAs
LoRA Experts Acquirement. Our proposed LoRA-Mixer framework is highly flexible and supports the integration of LoRA modules from diverse sources. In one common scenario, users may download pre-trained LoRA adapters from public repositories such as LoRAHub Huang et al. (2023), which currently hosts 196 high-quality LoRA modules across a wide range of domains. These can be directly composed using LoRA-Mixer with minimal additional data. Alternatively, users may independently train domain-specific LoRA modules tailored to their own datasets. For scenarios requiring joint training of multiple LoRA modules on a heterogeneous, labeled dataset, LoRA-Mixer further supports a hard-routing strategy. Specifically, we fix the routing module and apply a deterministic routing scheme based on known domain labels. Given a domain ID $d\in\{1,\dots,K\}$ associated with each training instance, all tokens within the sample are routed exclusively to expert $d$. This design enables efficient joint optimization while maintaining expert modularity. Collectively, these capabilities make LoRA-Mixer a versatile and scalable framework for composing heterogeneous LoRA modules. The overall architecture is illustrated in Figure 2.
@@ 3.3 Specialization Balance Loss for Routing Optimization
The next step is to optimize the expert routers. While previous research has introduced auxiliary loss functions to align the average gating score with expert utilization, thereby promoting load balancing, we observed that this approach overemphasizes consistency, resulting in an overly balanced expert distribution. In this case, all experts are forced to be used equally, regardless of the input semantics. This hinders effective routing and generally requires more training data.
We propose a new perspective: viewing the router as an information bottleneck that determines the degree to which semantic differences between tokens are preserved or compressed. From this perspective, the entropy of the routing distribution quantifies the uncertainty of the router. Therefore, load balancing and specialization selection appear to be in conflict. To achieve a balance between the two, ensuring balanced expert load and input-aware routing, we propose an improved optimization objective, called Routing Specialization Balance Loss (RSL). RSL addresses this issue by unifying the global budget and local selectivity into a single objective. Specifically, RSL enhances the auxiliary loss with an entropy regularization term, improving the accuracy of expert selection by suppressing overly flat distributions. Formally, let $\bar{p}_{i}$ denote the average soft routing score (across tokens) and $\bar{f}_{i}$ denote the normalized score assigned to the token of expert $i$ in the first $k$ routes. The RSL loss function is defined as:
@@
Where $\alpha$ controls the strength of the equilibrium consistency term, $\lambda$ is a small positive coefficient of the entropy regularization term:
@@
The design principles behind RSL are: 1) From an information bottleneck perspective, minimizing $\mathcal{H}(p(x))$ reduces token-conditional uncertainty under a fixed global load, directly promoting specialization without disrupting the balance. 2) The entropy term acts as a curvature provider in the routing simplex, resulting in strong convexity and well-conditioned optimization (see A.1); this is precisely where traditional auxiliary loss functions are weak. 3) The coefficient $\lambda$ becomes an interpretable knob that trades off global fairness and local specialization, rather than simply a regularization weight.
Specifically, the entropy term introduces token-specific entropy. We derive the entropy gradient with respect to the routing score $p_{i}(x)$:
$=-\log p_{i}(x)-1+\mu,$ (8)
where $\mu$ is the Lagrange multiplier due to the simplex constraint. Therefore, the total gradient of the RSL loss becomes:
This demonstrates that RSL introduces a token-level gradient signal via $\log p_{i}(x)$, unlike the auxiliary loss, which only propagates global gradients. Specifically, the $\log p_{i}(x)$ term provides a token-level signal that amplifies the input-aware variance in the expert’s choices, thereby counteracting the bias towards consistency in global balance. To quantify input-aware routing, we define:
We call routing token-aware if $\text{Var}(p(x))>\epsilon$ and $\epsilon>0$. The auxiliary loss function tends to reduce this variance (driving uniform routing), while RSL encourages high variance and peaked distributions that are consistent with the input semantics. In addition to the above proofs, we provide convergence analysis and generalization bound proofs in the Appendix A.1 A.2. The former introduces strong convexity on the product simplex and proves that the added entropy term can maintain stable optimization dynamics; the latter shows that entropy regularization reduces the assumed complexity of the router and improves robustness in low-data environments.
Routing Optimization. After LoRA is ready, we soft-train the router. To prevent expert knowledge from being contaminated, we introduce a regularization term to penalize deviations from the previously learned expert parameters. Let the first-stage parameters of expert $i$ be $\theta_{i}^{(0)}$ and the current parameters be $\theta_{i}$. We define the regularization term as:
@@
where $\mathcal{C}$ represents the set of constrained experts and $\beta$ controls the regularization strength. This regularization term constrains sensitive experts to maintain their original knowledge while allowing other experts to flexibly adjust, supporting multi-expert learning for complex tasks.
To ensure effective gradients and stable optimization during joint training, we employ soft expert fusion: the routers output a softmax score $\mathbf{p}_{b,t}\in\mathbb{R}^{K}$, achieving differentiable LoRA hybridization. While this approach provides stable gradients, combining it with an auxiliary loss results in equal activations among experts, inhibiting specialization. To address this, we introduce RSL, which promotes expert differentiation while maintaining load balancing and enables efficient reuse of LoRA with minimal data. The total loss during joint training is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\alpha\cdot\mathcal{L}_{\text{RSL}}+\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2},$ (12)
where $\mathcal{L}_{\text{task}}$ denotes the standard task loss (e.g., cross-entropy loss). $\mathcal{L}_{\text{RSL}}$ is the routing specialization balance loss, which is used to promote consistency between routing scores and actual token assignments. $\alpha$ is a weighting factor for RSL. $\mathcal{L}_{\text{preserve}}$ is the expert regularization loss described above, scaled by $\beta$. We provide an exploration of how to choose hyperparameters in Appendix A.8.
4 Experiment.
In this section, we conduct extensive experiments on 15 datasets across five domains (MedQA, commonsense reasoning, NLP, mathematics, and coding).